Grade 11Physical Chemistry

3.4gm of a metal requires 130.0ml of 0.5M HCl to dissolve it. What is equivalent weight of metal

Rahul
8 Years agoGrade 11

1 Answer

Vikas Kumar
8 Years ago
0.5=x/36.5 * 1000/130 (Molarity formula)
x=2.375gm
Eq wt of HCl = 36.5 as replaceable H is only 1
So, by W1/E1=W2/E2
Eq wt of metal = 52.2